Your parents will retire in 29 years. They currently have $290,000 saved, and they think they will need $2,100,000 at retirement. What annual interest rate must they earn to reach their goal, assuming they don't save any additional funds? Round your answer to two decimal places.
Answer #1

Ans 7.07%

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
2100000 / 290000 = (1 + r%)^29
7.24137931034483 = (1 + r%)^29
r = 7.07%